40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 12 17:22:07 2013, Steve, Update, General, rack power supplies checked 8x
    Reply  Fri Dec 13 16:34:20 2013, Steve, Update, VAC, PS will be replaced next week hotPSvacRack.jpgfan1X8.jpg
       Reply  Fri Dec 20 20:04:01 2013, Steve, Update, VAC, power supply replaced with a short vent 4hrPumpdown.pngpumpdownAfterHickup.pngPSpumpdown.png
          Reply  Fri Dec 20 23:00:41 2013, Koji, Update, VAC, power supply replaced with a short vent 
             Reply  Sat Dec 21 10:53:35 2013, Steve, Update, VAC, power supply replaced with a short vent-pumpdown completed 20hrsVacNormal.png
          Reply  Fri Jan 3 11:18:41 2014, Steve, Summary, VAC, power supply replaced with a short vent rebootVENT.png
    Reply  Thu Feb 22 10:49:11 2018, Steve, Update, Electronics, rack power supplies checked 8x
Message ID: 9516     Entry time: Fri Jan 3 11:18:41 2014     In reply to: 9506
Author: Steve 
Type: Summary 
Category: VAC 
Subject: power supply replaced with a short vent 

Quote:

Quote:

 

 The temperature went down to room temp with temporary fan in the back. Voltage and current are stable.

Regardless, it will be replaced early next week.

Koji, Steve

 It was a bad experience again with our vacuum system.  The valves went crazy as we rebooted the computer. This was required for the swap in of a good 24V power supply.

The IFO was vented to 27 Torr through the annuloses, VA6, V7, Maglev,VM2 and VM1 (VC2 was open too)

I just opened the PSL shutter after a 4 hours pumpdown.

Condition: annuloses are not pumped, the IFO and the RGA are pumped as Atm2 shows

I will be here tomorrow morning to switch over to vacuum normal. 

More details later

 

 

 Events of the power supply swap:

1, Tested 24V DC ps from Todd

2, Closed V1, VM1 and all annulos valves to create safety net for the reboot. Turbo pumps left on running.

3, Turned computer off

4, Swap power supplies and turned it on

5, Turning the power on of c1vac2 created caos switching of valves. This resulted in a air vent as shown below.

6, VM1 was jammed and it was unable to close. The IOO beam shutter closed and the IFO was venting with air for a few minutes. Maglev did an emergency shut down. TP2's V4 and TP3' V5 closed. RP1 and RP3                           roughing         pumps turned on, their hose was not connected as usual. The RGA shut down to protect itself.

7, Closed annulos valves, stoped the vent at P1 27 torr as the vacuum control was  manually recovered

8, The Maglev and the annuloses were roughed out 500 mtorr . The Maglev was restarted.

9, The IFO pump down followed std procedure from 27 torr. VM1 was moving again as the pressure differential was removed from it.

 

 Remember: next time at atm .....rough down the cryo volume from 27 torr !

Attachment 1: rebootVENT.png  351 kB  | Hide | Hide all
rebootVENT.png
ELOG V3.1.3-